OpenStudy (anonymous):

what is the area of a circle with a radius of 15in to the nearest tenth

OpenStudy (anonymous):

2 * pi * r = 2 * 22/7 * 15 = 94.3 square inches.

OpenStudy (anonymous):

medals?

OpenStudy (anonymous):

????????????????????????????????????????

OpenStudy (anonymous):

M-E-D-A-L

OpenStudy (anonymous):

i'd leave it in pi, as you rounded the actual answer !! XD

OpenStudy (anonymous):

and A = PI R^2...

OpenStudy (anonymous):

so your answer should be 225 pi

OpenStudy (eyust707):

no soup fo you!